a penn state professor drives from state college to tyrone, 20 miles away, at 40 mph. He stops for 20 minutes in Tyrone for coffee, then drives to Altoona (another 20 miles away) at 60 mph. Assuming State College, Tyrone and Altoona to be more or less on a straight line, what is the magnitude of the professor's average velocity for the entire trip?

Please show work so that I can understand how to solve the problem on my own for future problems
1
Expert's answer
2018-09-12T09:47:08-0400
We need to find average speed:
V=D/T.
t_2=20 min=20/60 h.
We have 3 parts of journey. The total time will be:
T=t_1+t_2+t_3=d_1/v_1 +t_2+d_3/v_3 =20/40+20/60+20/60=7/6 h.
D=d_1+d_2+d_3=20+0+20=40 miles.
Thus,
V=40/(7/6)≈34 mph.